The opening balance in the Provision for Doubtful Debts was Rs.705. During the year, the direct write-offs amounted to Rs.30 and at the year end the Debtors account showed a balance of Rs.15,000 (after writing-off Rs.30). The charge to profit and Loss account for bad debt expense for the year including the direct write-off amounted to Rs.200. Calculate the percentage provision that had been made for doubtful debts at the year end
1
Expert's answer
2016-10-06T12:55:04-0400
Opening Balance in the Provision = Rs. 705

Less Bad Debts of 2016 = Rs. 30

Balance amount left in the Provision = Rs. 675

Provision required for 2017 (15,000 X ?/100) = Rs. ?

Excess amount in the Provision to be credited to P & L A/c = Rs. 200

So, Provision required for 2017 = 675 – 200 = 475
15,000 X ? / 100 = 475
Percentage provision = 475 / 15,000 X 100 = 3.167% ≈ 3.2 %

Percentage provision doubtful debts is 3.2 %.